Complete this question by entering your answers In the tabs below. Req a and b| Req c Req d to h d. A single person with taxable income of $80,136 e. A single person with taxable income of $313,471 f. A head of household with taxable income of $97,292. g. A qualifying widow with taxable income of $25,856. h. A married couple filing jointly with taxable income of $11,916. tax rate 410 % 270% 24 00% 350 % 240 0% 1200% 1000 g- 1660 % Req c K Prev 3 of 4 Next > 回 Do al 8 3 4 5 6 8

%201259917088/2018-tax-rate-schedule pdf endix F 2018 Federal Income Tax Information 2018 Federal Tax Rate Schedules Schedule X-Single If taxable income is over: But not over: The tax is: S0 $9,525 $38,700 $82.500 $157,500 $200.000 $500,000 $9,525 $38,700 10% of taxable income $952.50 plus 1 2% of the excess over S 9.525 $4,453.50 plus 22% of the excess over $38,700 $14,089 50 plus 24% of the excess over S82.500 $32,089.50 plus 32% of the excess over $ 1 57,500 $45,689 50 plus 35% of the excess over $200,000 $150.689 50 plus 37% of the excess over S500000 $157,500 $200,000 $500,000 Schedule Y1-Married Filing Jointly or Qualifying Widow(er) If taxable The tax is: But not over: $19,050 $77,400 $165.000 5 315,000 $400.000 is over: s0 $19.050 $77,400 $165.000 $315.000o $400,000 10% of taxable income $1 .905.00 plus 1 2% of the excess over $ 19.050 sa 907 00 plus 22% of the excess over $77.400 $28.179 00 plus 24% of the excess over $ 165.000 $64. I 79.00 plus 32% of the excess over $315,000 $91 . 3 79 00 plus 35% of the oxcess ouer S400.000 $161 379.00 plus 37%ofthe excess over 5600,000 $600 回

$600,000 $16 1 ,379.00 plus 37% of the excess over $600,000 Schedule Y-2-Married Filing Separately H taxable Income is over: But not over: The tax is: $0 $9,525 $38,700 $82,500 $157,500 $200,000 $300,000 $9,525 $38,700 $82,500 $157,500 $200,000 $300,000 10% of taxable income $952.50 plus 1 2% of the excess over $9,525 $4,453.50 plus 22% of the excess over $38,700 $14,089 50 plus 24% ofthe excess over $82,500 $32,089.50 plus 32% of the excess over $157,500 $45,689.50 plus 35% of the excess over $200,000 $80,689.50 plus 37% of the excess over $300,000 Schedule Z-Head of Household If taxable income is over: The tax is: But not over: $0 $13,600 $51,800 $82,500 $157,500 $200,000 $500,000 $13,600 $51 ,800 $82.500 $ 1 57,500 $200,000 $500,000 10% of taxable income $ 1 .360 00 plus 1 2% of the excess over $ 13,600 $5,944.00 plus 22% of the excess over $51,800 $ 1 2.698.00 plus 24% of the excess over $82,500 $30,698 00 plus 32% of the excess over $157,500 $44,298 00 plus 35% of the excess over $200,000 $1 49.298 00 plus 37% of the excess over $500,000 F-1
Answer #1

d.Taxable income of an individual between $38700 and $82500 will be taxed at 22%.In given case, taxable income is $80136. It will come in third tax bracket..In given case, taxable income is $80136. It will come in third tax bracket.The tax rate corresponding to the bracket in which taxable income falls is the marginal tax rate.

This individual's taxable income falls in 3rd tax bracket and tax rate corresponding to second tax bracket is 22%.

So,

This individual's marginal tax rate is 22 %

Tax payable =4453.50+(80136-38700)*22%

=13,569.42

Average tax rate =(Tax payable /Taxable income) *100

=(13,569.42/80136)*100

=16.93%

e. Taxable income of an individual between $200000 and $500000 will be taxed at 35%.

Marginal tax rate =35%

Tax payable =45689.50+(313471-200000)*35%

=85,404.35

Average tax rate =(Tax payable /Taxable income) *100

=(85,404.35/313471)*100

=27.24%

f. Taxable income of a head of household between $82500 and $157500 will be taxed at 24%.

Marginal tax rate =24%

Tax payable =12698+(97292-82500)*24%

=16,248.08

Average tax rate =(Tax payable /Taxable income) *100

=(16,248.08/97292)*100

=16.70%

g. Taxable income of a qualifying widow between $19050and $77400 will be taxed at 12%.

Marginal tax rate =12%

Tax payable =1905+(25856-19050)*12%

=2,721.72

Average tax rate =(Tax payable /Taxable income) *100

=(2,721.72/25856)*100

=10.53%

h. Taxable income of married couple between $0 and $19050 will be taxed at 10%.

Marginal tax rate =10%

Tax payable =11916*10%

=1191.6

Average tax rate =(1191. 6/11916)*100

=10%

IN CASE OF ANY DOUBTS OR CORRECTIONS FEEL FREE TO COMMENT BELOW

PLEASE RATE MY ANSWER BY HITTING ??

THANK YOU